Reports 1-1 of 1 Clear search Modify search
MIF (ASC)
dan.chen - 17:14 Monday 02 February 2026 (36259) Print this report
Third trial of high-bandwidth ASC: DHARD pitch

[Komori, Dan]

We adjusted the DHARD P filter "HBtest" to increase the UGF of the DHARD P loop.

We

  • adjusted the notch filter freq (it was at 7.5Hz but we moved it to 7.35Hz based on the TF meas data.)
  • added another notch filter at 9.2Hz (after the work, we found it is near by the freq of PRM ADS, which was ON during the meas. So it might be not nessesary.)

After these change, we could raise the gain of DHARD P loop to -3.5 which gave the 2Hz  UGF.

Meas file: /users/Commissioning/data/ASC/2026/0202/DHARD_P_newact_OLTF.xml

Images attached to this report
Comments to this report:
dan.chen - 17:46 Monday 02 February 2026 (36260) Print this report

[Tanaka, Komori, Dan]

We adjuseted the ASC OUTMTRX for {D,C}{HARD,SOFT} to reduce the couplings.

(After some trial with different methods to estimated the appropriate values) we evaluate the values with the following method:

  1. Requested LSC, ASC to be down
  2. Keep the suspension to be LOCK_AQUISITION state
  3. Turned OFF MN MNOLDAMP of all type A suspension by turning ON FM8=null filter of them.
  4. Cut the path of TM by change the gain of TM LOCK from 1 to 0 for all type A suspensions.
  5. Injected a sin (freq = 0.05Hz) into K1:ASC-DHARD_P_SM_EXC and measured TM oplev signals.

We used the measured values to calculate the OUTMTRX values and updated them and measure it again to confirm that the oplev values gave almost same values.
(We have NOT confirmed that the couplings are redueced.)

Attached figures are the measured TF from K1:ASC-DHARD_P_SM_EXC to oplevs and the OUTMTRX after the adjustment.

We confirmed that the IFO can reach PRFPMI_RF_LOCKED state at the end with the new OUTMTRX values.

Images attached to this comment
Search Help
×

Warning

×